geerky42 (geerky42):

Find the \(\large 10^{-8}\) place of \(\sin (2)\).

OpenStudy (anonymous):

What do you mean by place?

geerky42 (geerky42):

Value of a digit depends on its place.

OpenStudy (anonymous):

Is this from taylor series?

geerky42 (geerky42):

Original question: \[f(x) = (x-2) - \dfrac{(x-2)^3}{3!} + \dfrac{(x-2)^5}{5!} - \dfrac{(x-2)^7}{7!} \cdots\] Find the \(10^{-8}\) place of \(f(4)\).

geerky42 (geerky42):

Yes it appears so. @incomplte